given f(x)=1/x and g(x)=2/x^2 find f+g(-3)
a.-5/9
b.-1/9
c.1/2
d. 1/12
Thanks

Answer by Edwin McCravy(20054)


given f(x)=1/x and g(x)=2/x^2 find f+g(-3) 
a.-5/9
b.-1/9
c.1/2
d. 1/12 
Thanks

That just means to add f(-3) and g(-3)

 f(x) = 1/x
f(-3) = 1/(-3)
f(-3) = -1/3

 g(x) = 2/x²
g(-3) = 2/(-3)²
g(-3) = 2/9

Therefore (f+g)(-3) = f(-3)+g(-3) 
                    = -1/3 + 2/9 
                    = -3/9 + 2/9 
                    = -1/9  (choice b)

Edwin